OpenStudy (anonymous):

Trig exact value on interval [0,2pi)

OpenStudy (anonymous):

tan^-1 sqrt3/3

OpenStudy (jdoe0001):

I'd think you'd just use the calculator for that one same as you'd on the last one as well

OpenStudy (anonymous):

Don't allow us to use one. My main thing is determine what quadrants to use.

OpenStudy (jdoe0001):

I assume you have a Unit Circle with values of common angles? at least with the cosine and sine, right?

OpenStudy (anonymous):

Yes I do.

OpenStudy (jdoe0001):

bear in mind that that \(\bf tan^{-1}\left( \cfrac{\sqrt{3}}{3} \right)=\theta\qquad tan(\theta)=\cfrac{\sqrt{3}}{3}\leftarrow \cfrac{sin(\theta)}{cos(\theta)} \\ \quad \\ also\qquad \cfrac{sin(\theta)}{cos(\theta)}\to \cfrac{1}{\sqrt{3}}\implies \cfrac{1}{\sqrt{3}}\cdot \cfrac{\sqrt{3}}{\sqrt{3}}\implies \cfrac{\sqrt{3}}{3}\)

OpenStudy (jdoe0001):

so.. whatever sine/cosine were, their division ended up as \(\bf \cfrac{1}{\sqrt{3}}\)

OpenStudy (jdoe0001):

that usually means that the cosine is the one with the radical the one I can see right off is \(\bf \cfrac{\pi }{6}\)

OpenStudy (anonymous):

Alright I think I'm getting it now. Thank you so much for the help!